OpenStudy (anonymous):

How do I find sin(170) degrees without using a calculator?

OpenStudy (anonymous):

there is a chart that you can consult, but other than that, not possible

OpenStudy (anonymous):

0.346649455

OpenStudy (anonymous):

In the absence of either a table or a chart you could estimate it. It will be the sin (180 -10). Now since sin 90 = 1 and sin 180 = 0 then naively using linear extrapolation would imply sin 10 = 1/9 or roughly 0.11 so that would give sin 170 = 0.11 - it's actually 0.17 so you can see the shortcoming of assuming linear but like I say it's an estimate not a guess. See if anyone can do better.

OpenStudy (anonymous):

If you recall that sin 30 is special ie it's 0.5 then sin 10 = 0.5/3 roughly = 0.17 and obviously sin 170 = sin 10 so there you go provided you recall that sin 30 = 0.5